Write an equation of the line in slope-intercept form with slope 1/3 that crosses the y-axis at (0, -2)
Answer by venugopalramana(3286)

Write an equation of the line in slope-intercept form with slope 1/3 that crosses the y-axis at (0, -2)
SLOPE INTERCEPT IS
Y=MX+C
M=1/3 AND Y-INTERCEPT = -2 ...SO EQN.IS....
Y=(X/3)-2
3Y=X-6